### v1

# Rethinking Consistent Multi-Label Classification under Inexact Supervision

Abstract
Partial multi-label learning and complementary multi-label learning are two popular weakly supervised multi-label classification paradigms that aim to alleviate the high annotation costs of collecting precisely annotated multi-label data. In partial multi-label learning, each instance is annotated with a candidate label set, among which only some labels are relevant; in complementary multi-label learning, each instance is annotated with complementary labels indicating the classes to which the instance does not belong. Existing consistent approaches for the two paradigms either require accurate estimation of the generation process of candidate or complementary labels or assume a uniform distribution to eliminate the estimation problem. However, both conditions are usually difficult to satisfy in real-world scenarios. In this paper, we propose consistent approaches that do not rely on the aforementioned conditions to handle both problems in a unified way. Specifically, we propose two unbiased risk estimators based on first- and second-order strategies. Theoretically, we prove consistency w.r.t. two widely used multi-label classification evaluation metrics and derive convergence rates for the estimation errors of the proposed risk estimators. Empirically, extensive experimental results validate the effectiveness of our proposed approaches against state-of-the-art methods.
1 Introduction
In multi-label classification (MLC), each instance is associated with multiple relevant labels simultaneously (Zhang and Zhou, 2014; Liu et al., 2022b). The goal of MLC is to induce a multi-label classifier that can assign multiple relevant labels to unseen instances. MLC is more practical and useful than single-label classification, as real-world objects often appear together in a single scene. The ability to handle complex semantic information has led to the widespread use of MLC in many real-world applications, including multimedia content annotation (Cabral et al., 2011), text classification (Rubin et al., 2012; Liu et al., 2017), and music emotion analysis (Wu et al., 2014). However, annotating multi-label training data is more expensive and demanding than annotating single-label data. This is because each instance can be associated with an unknown number of relevant labels (Durand et al., 2019; Cole et al., 2021; Xie et al., 2023), making it difficult to collect a large-scale multi-label dataset with precise annotations.
Figure 1: A multi-label image with inexact annotations. Source: Paul Cézanne, Still Life, Jug and Fruit on a Table (1894), public domain.
To address this, learning from weak supervision has become a prevailing way to mitigate the bottleneck of annotation cost for MLC (Sugiyama et al., 2022). Among them, partial multi-label learning (PML) and complementary multi-label learning (CML) have become two popular MLC paradigms. In PML, each instance is annotated with a candidate label set, among which only some labels are relevant but inaccessible to the learning algorithm (Xie and Huang, 2018; Sun et al., 2019; Gong et al., 2021). In CML, each instance is annotated with complementary labels, which indicate the classes to which the instance does not belong (Gao et al., 2023). Given that all relevant labels are included in the candidate label set, non-candidate labels contain no relevant labels and can be considered complementary labels, and vice versa. This suggests that the two problems are mathematically equivalent. Therefore, in this paper, we treat them as MLC under inexact supervision in a unified way. Figure 1 shows an example image annotated with inexact annotations. The label space contains ten labels in total. The candidate label set consists of four relevant labels {apple, plate, table, jug} and two false-positive ones {grapes, pear}. By excluding the candidate labels from the label space, the remaining four labels are {banana, cup, knife, flower}, which can be considered complementary labels. PML and CML do not require precise determination of all relevant labels during annotation, which demonstrates their great potential for alleviating annotation challenges in MLC.
In this paper, we investigate consistent approaches for MLC under inexact supervision. Here, consistency means that classifiers learned with inexact supervision are theoretically guaranteed to converge to the optimal classifiers when infinitely many training samples are provided (Wang et al., 2024). The remedy began with Xie and Huang (2023), which treated PML as a special case of MLC with class-conditional label noise (Li et al., 2022; Xia et al., 2023), where irrelevant labels could flip to relevant labels but not vice versa. However, the flipping rate for each class is unknown and must be estimated using anchor points, i.e., instances belonging to a specific class with probability one (Liu and Tao, 2015; Xie and Huang, 2023). Similar to PML, CML assumes that complementary labels are generated by a certain flipping process (Yu et al., 2018b). Gao et al. (2023) proposed the uniform distribution assumption that a label outside the relevant label set is sampled uniformly to be the CL. Then, Gao et al. (2024) generalized the data generation process with a transition matrix, but estimating the data generation process is still necessary. Recently, Gao et al. (2025) extended the uniform distribution assumption to handle multiple complementary labels.
In summary, all existing consistent PML and CML approaches either estimate the generation process of the candidate label set or complementary labels, or adopt the uniform distribution assumption to eliminate the estimation problem. However, both conditions are difficult to satisfy in real-world scenarios. On the one hand, estimating the flipping rate heavily relies on accurate estimation of noisy class posterior probabilities of anchor points (Xia et al., 2019; Yao et al., 2020; Lin et al., 2023). However, estimating noisy class posterior probabilities is more difficult because their entropy is usually higher than that of clean labels (Langford, 2005). This difficulty is further amplified when using deep neural networks, where the over-confidence phenomenon typically occurs (Zhang et al., 2021; Wei et al., 2022). The model outputs of deep neural networks are usually one-hot encoded, which means they cannot yield reliable probabilistic outputs (Guo et al., 2017). On the other hand, the uniform distribution assumption treats different candidate label sets indiscriminately, which is too simple to be truly in accordance with imbalanced classes in real-world scenarios (Wang et al., 2025). Additionally, many approaches model different labels independently and directly ignore label correlations existing in multi-label data (Gao et al., 2025). This prevents them from exploiting the rich semantic relationships of label correlations (Zhu et al., 2017; Mao et al., 2023).
Table 1: Comparison of COMES with existing consistent PML and CML approaches.
Approach Uniform distribution assumption-free Generation process estimation unnecessary Label correlation-aware Multiple complementary labels
CCMN (Xie and Huang, 2023)
CTL (Gao et al., 2023)
MLCL (Gao et al., 2024)
GDF (Gao et al., 2025)
COMES-HL (Ours)
COMES-RL (Ours)
To this end, we propose a novel framework named COMES, i.e., COnsistent Multi-label classification under inExact Supervision. Based on a data generation process that does not use transition matrices, we introduce two instantiations with unbiased risk estimators w.r.t. the Hamming loss and ranking loss, respectively. Table 1 compares our approach with existing consistent PML and CML approaches. Our contributions are summarized as follows:
  • We propose a general and consistent framework for multi-label classification under inexact supervision that neither requires estimating the generation process of candidate or complementary labels nor relies on the uniform-distribution assumption.
  • We introduce risk-correction approaches to improve the generalization performance of the proposed unbiased risk estimators. We further prove consistency w.r.t. two widely used metrics and derive convergence rates of estimation errors for the proposed risk estimators.
  • Our proposed approaches outperform state-of-the-art baselines on both real-world and synthetic PML and CML datasets with different label generation processes.
2 Preliminaries
In this section, we introduce the background of MLC and MLC under inexact supervision.
2.1 Multi-Label Classification
Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ denote the $d$-dimensional feature space and $\mathcal{Y}=\left\{1,2,\ldots,q\right\}$ the label space consisting of $q$ class labels. A multi-label example is denoted as $(\bm{x},Y)$, where $\bm{x}\in\mathcal{X}$ is a feature vector and $Y\subseteq\mathcal{Y}$ is the set of relevant labels associated with $\bm{x}$. For ease of notation, we introduce $\bm{y}=[y_{1},y_{2},\ldots,y_{q}]\in\left\{0,1\right\}^{q}$ to denote the vector representation of $Y$, where $y_{j}=1$ if $j\in Y$ and $y_{j}=0$ otherwise. Let $p(\bm{x},Y)$ denote the joint density of $\bm{x}$ and $Y$. Let $p(\bm{x})$ denote the marginal density, and $\pi_{j}=p(y_{j}=1)$ the prior of the $j$-th class. The task of MLC is to learn a prediction function $\bm{f}:\mathcal{X}\mapsto 2^{\mathcal{Y}}$. We use $f_{j}$ to denote the $j$-th entry of $\bm{f}$, where $f_{j}(\bm{x})=1$ indicates that the model predicts class $j$ to be relevant to $\bm{x}$ and $f_{j}(\bm{x})=0$ otherwise. Since learning $\bm{f}$ directly is often difficult, we use a real-valued decision function $\bm{g}:\mathcal{X}\mapsto\mathbb{R}^{q}$ to represent the model output. The prediction function $\bm{f}$ can be derived by thresholding $\bm{g}$. We use $g_{j}$ to denote the $j$-th entry of $\bm{g}$, which indicates the model output for class $j$.
Many evaluation metrics have been developed to calculate the difference between model predictions and true labels to evaluate the performance of multi-label classifiers (Zhang and Zhou, 2014; Wu and Zhou, 2017). In this paper, we focus primarily on the Hamming loss and ranking loss, the two most common metrics in the literature. [^1] We will address the use of other metrics in future work. Specifically, the Hamming loss calculates the fraction of misclassified instance-label pairs, and the risk of $\bm{f}$ w.r.t. the Hamming loss is
$R_{\mathrm{H}}^{\mathrm{0-1}}(\bm{f})=\mathbb{E}_{p(\bm{x},Y)}\left[\frac{1}{q}\sum\nolimits_{j=1}^{q}\mathbb{I}\left(f_{j}(\bm{x})\neq y_{j}\right)\right].$ (1)
Here, $\mathbb{I}$ denotes the indicator function that returns 1 if the predicate holds; otherwise, $\mathbb{I}$ returns 0. Since optimizing the 0-1 loss is difficult, a surrogate loss function $\ell$ is often adopted. The $\ell$-risk w.r.t. the Hamming loss is
$R_{\mathrm{H}}^{\ell}(\bm{g})=\mathbb{E}_{p(\bm{x},Y)}\left[\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),y_{j}\right)\right],$ (2)
where $\ell$ is a non-negative binary loss function, such as the binary cross-entropy loss. It is important to note that the Hamming loss only considers first-order model predictions and cannot account for label correlations. The ranking loss explicitly considers the ordering relationship between model outputs for a pair of labels. Specifically, the risk of $\bm{f}$ w.r.t. the ranking loss is [^2] To facilitate the analysis in this paper, we consider the coefficients of the losses for different label pairs to be 1 (Gao and Zhou, 2013; Xie and Huang, 2021, 2023).
$R_{\mathrm{R}}^{\mathrm{0-1}}(\bm{f})=\mathbb{E}_{p(\bm{x},Y)}$ $\left[\sum\nolimits_{1\leq j<k\leq q}\mathbb{I}(y_{j}<y_{k})\left(\mathbb{I}\left(f_{j}(\bm{x})>f_{k}(\bm{x})\right)+\frac{1}{2}\mathbb{I}\left(f_{j}(\bm{x})=f_{k}(\bm{x})\right)\right)\right$.
$\left.+\mathbb{I}(y_{j}>y_{k})\left(\mathbb{I}\left(f_{j}(\bm{x})<f_{k}(\bm{x})\right)+\frac{1}{2}\mathbb{I}\left(f_{j}(\bm{x})=f_{k}(\bm{x})\right)\right)\right].$ (3)
Similarly, when using a surrogate loss function $\ell$ to replace the 0-1 loss, the $\ell$-risk w.r.t. the ranking loss is
$R_{\mathrm{R}}^{\ell}(\bm{g})=\mathbb{E}_{p(\bm{x},Y)}\left[\sum\nolimits_{1\leq j<k\leq q}\mathbb{I}(y_{j}\neq y_{k})\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),\frac{y_{j}-y_{k}+1}{2}\right)\right].$ (4)
Notably, minimizing the Hamming loss does not consider label correlations and can be considered a first-order strategy. In contrast, minimizing the ranking loss considers label-ranking relationships and can be considered a second-order strategy.
2.2 Multi-Label Classification under Inexact Supervision
In PML, each example is denoted as $\left(\bm{x},S\right)$, where $S$ is the candidate label set associated with $\bm{x}$. The basic assumption of PML is that all relevant labels are contained within the candidate label set, i.e., $Y\subseteq S$. Let $\S\bar{S}=\mathcal{Y}\backslash$ S denote the absolute complement of $S$. Since $\bar{S}\bigcap Y=\emptyset$, $\bar{S}$ can be regarded as the set of complementary labels associated with $\bm{x}$. Therefore, PML and CML are mathematically equivalent, as partial multi-label data can equivalently be transformed into complementary multi-label data and vice versa. Without loss of generality, this paper mainly considers partial multi-label data. For ease of notation, we use $\bm{s}=[s_{1},s_{2},\ldots,s_{q}]$ to denote the vector representation of $S$. Here, $s_{j}=1$ indicates that the $j$-th class label is a candidate label of $\bm{x}$, and $s_{j}=0$ otherwise. Let $p(\bm{x},S)$ denote the joint density of $\bm{x}$ and the candidate label set $S$. The goal of PML or CML is to learn a prediction function $\bm{f}:\mathcal{X}\mapsto 2^{\mathcal{Y}}$ that can assign relevant labels to unseen instances based on a training set $\mathcal{D}=\left\{\left(\bm{x}_{i},S_{i}\right)\right\}_{i=1}^{n}$ sampled i.i.d. from $p(\bm{x},S)$.
3 Methodology
In this section, we first introduce our data generation process. Then, we present the first- and second-order strategies for handling the PML problem and their respective theoretical analyses.
3.1 Data Generation Process
In this paper, we assume that the candidate labels are generated by querying whether each instance is irrelevant to a class in turn. Specifically, if the $j$-th class is irrelevant to $\bm{x}$, we assume that the $j$-th class label is assigned as a non-candidate label to $\bm{x}$ with a constant probability $p_{j}$, i.e., $p\left(j\notin S|\bm{x},j\notin Y\right)=p_{j}$. Otherwise, if the $j$-th class is relevant to $\bm{x}$, we consider it as a candidate label. The candidate label set can then be obtained by excluding the non-candidate labels from the label space. Notably, all relevant labels are included in the candidate label set, as well as some irrelevant labels. This data generation process coincides well with the annotation process of candidate labels. For example, when asking annotators to provide candidate labels for an image dataset, we can show them an image and a class label and ask them to determine whether the image is irrelevant to that class. This is often an easier question to answer than directly asking all relevant labels, since it is less demanding to exclude some obviously irrelevant labels. If so, we assume that the image will be annotated with this label as a non-candidate label with a constant probability. Based on this data generation process, we have the following lemma.
Lemma 1.
Assume that $p(s_{j}=0|\bm{x},y_{j}=0)=p_{j}$, where $p_{j}$ is a constant. Then, we have $p(\bm{x}|s_{j}=0)=p(\bm{x}|y_{j}=0)$.
The proof can be found in Appendix B.1. According to Lemma 1, the conditional density of instances where the $j$-th class is considered a non-candidate label is equivalent to the conditional density of instances where the $j$-th class is irrelevant. Notably, our data distribution assumption differs from both the uniform distribution assumption and the use of a transition matrix to flip the labels. Since the conditional probabilities of different candidate label sets can be different, our setting is more general than the uniform distribution assumption (Gao et al., 2023, 2025).
3.2 First-Order Strategy
A common strategy used in MLC is to decompose the problem into a number of binary classification problems by ignoring label correlations. This goal can be achieved by minimizing the $\ell$-risk w.r.t. the Hamming loss in Eq. (2). We show that the $\ell$-risk w.r.t. the Hamming loss can be equivalently expressed with partial multi-label data.
Theorem 1.
By the assumption in Lemma 1, the $\ell$-risk w.r.t. the Hamming loss in Eq. (2) can be equivalently expressed as
$R_{\mathrm{H}}^{\ell}(\bm{g})=$ $\mathbb{E}_{p(\bm{x})}\left[\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),1\right)\right]$
$+\sum\nolimits_{j=1}^{q}\mathbb{E}_{p(\bm{x}|s_{j}=0)}\left[\frac{1-\pi_{j}}{q}\left(\ell\left(g_{j}\left(\bm{x}\right),0\right)-\ell\left(g_{j}\left(\bm{x}\right),1\right)\right)\right].$ (5)
The proof can be found in Appendix B.2. Theorem 1 shows that the $\ell$-risk w.r.t. the Hamming loss can be expressed as the expectation w.r.t. the marginal and conditional densities where the $j$-th class label is not considered as a candidate label. Since Eq. (1) cannot be calculated directly, we perform empirical risk minimization (ERM) by approximating Eq. (1) using datasets $\mathcal{D}_{\mathrm{U}}$ and $\mathcal{D}_{j}(j\in\mathcal{Y})$ sampled from densities $p(\bm{x})$ and $p(\bm{x}|s_{j}=0)$, respectively. In this paper, we consider generating these datasets by duplicating instances from $\mathcal{D}$. Specifically, we first treat the duplicated instances of $\mathcal{D}$ as unlabeled data sampled from $p(\bm{x})$ and add them to $\mathcal{D}_{\mathrm{U}}$. Then, if an instance does not treat the $j$-th class label as a candidate label, we treat its duplicated instance as being sampled from $p(\bm{x}|s_{j}=0)$ and add it to $\mathcal{D}_{j}$. These processes can be expressed as follows:
$\mathcal{D}_{\mathrm{U}}=\left\{\bm{x}^{\mathrm{U}}_{i}\right\}_{i=1}^{n}=\left\{\bm{x}_{i}|(\bm{x}_{i},S_{i})\in\mathcal{D}\right\},\quad\mathcal{D}_{j}=\left\{\bm{x}^{j}_{i}\right\}_{i=1}^{n_{j}}=\left\{\bm{x}_{i}|(\bm{x}_{i},S_{i})\in\mathcal{D},j\notin S_{i}\right\},j\in\mathcal{Y}.$ (6)
Then, an unbiased risk estimator can be derived to approximate Eq. (1) using datasets $\mathcal{D}_{\mathrm{U}}$ and $\mathcal{D}_{j}$:
$\hat{R}_{\mathrm{H}}^{\ell}(\bm{g})=$ $\frac{1}{nq}\sum\nolimits_{i=1}^{n}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)$
$+\sum\nolimits_{j=1}^{q}\frac{1-\pi_{j}}{qn_{j}}\sum\nolimits_{i=1}^{n_{j}}\left(\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),0\right)-\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\right).$ (7)
When deep neural networks are used, the negative terms in the loss function can often lead to overfitting issues (Kiryo et al., 2017; Sugiyama et al., 2022). Therefore, we use an absolute value function to wrap each potentially negative term Lu et al. (2020); Wang et al. (2023). The corrected risk estimator is defined as
$\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})=$ $\frac{1}{q}\sum\nolimits_{j=1}^{q}\left|\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\right|$
$+\sum\nolimits_{j=1}^{q}\frac{1-\pi_{j}}{qn_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),0\right).$ (8)
Notably, our framework is very flexible so that the minimizer can be obtained using any network architecture and stochastic optimizer. The algorithmic details are summarized in Algorithm 1. The class prior $\pi_{j}$ can be estimated by using off-the-shelf class prior estimation approaches only using candidate labels (see Appendix A.2).
We establish the consistency and estimation error bounds for the risk estimator proposed in Eq. (3.2). First, we demonstrate that the corrected risk estimator in Eq. (3.2) is biased yet consistent w.r.t. the $\ell$-risk w.r.t. the Hamming loss in Eq. (2). The following theorem holds.
Theorem 2.
Assume that there exists a constant $C_{\mathcal{G}}$ such that $\sup_{g_{j}\in\mathcal{G}}\|g\|_{\infty}\leq C_{\mathcal{G}}$ and a constant $C_{\ell}$ such that $\sup_{|z|\leq C_{\mathcal{G}}}\ell(z,y)\leq C_{\ell}$, where $\mathcal{G}$ is the model class. Since $\ell$ is non-negative, we assume that there exists a non-negative constant $\alpha$ such that $\forall j\in\mathcal{Y},\pi_{j}\mathbb{E}_{p(\bm{x}|y_{j}=1)}\left[\ell\left(g_{j}\left(\bm{x}\right),1\right)\right]\geq\alpha$. Then, the bias of the expectation of the corrected risk estimator w.r.t. the $\ell$-risk w.r.t. the Hamming loss has the following lower and upper bounds:
$0\leq\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-R_{\mathrm{H}}^{\ell}(\bm{g})\leq\frac{1}{q}\sum\nolimits_{j=1}^{q}(4-2\pi_{j})C_{\ell}\Delta_{j},$ (9)
where $\Delta_{j}=\exp\left(-2\alpha^{2}/\left(C_{\ell}^{2}/n+(1-\pi_{j})^{2}C_{\ell}^{2}/n_{j}\right)\right)$. Furthermore, for any $\delta>0$, the following inequality holds with probability at least $1-\delta$:
$\left|\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})-R_{\mathrm{H}}^{\ell}(\bm{g})\right|\leq\frac{1}{q}\sum\nolimits_{j=1}^{q}\left((4-2\pi_{j})C_{\ell}\Delta_{j}+\frac{(2-2\pi_{j})C_{\ell}}{q}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n_{j}}}\right)+C_{\ell}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n}}.$ (10)
The proof can be found in Appendix B.3. Notably, the bias of the corrected risk estimator from the original $\ell$-risk exists since it is lower bounded by zero. However, as $n\rightarrow\infty$, we have that $\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\rightarrow R_{\mathrm{H}}^{\ell}(\bm{g})$, meaning that it is still consistent.
Let $\tilde{\bm{g}}_{\mathrm{H}}=\mathop{\arg\min_{\{g_{j}\}\subseteq\mathcal{G}}}\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})$ and $\bm{g}^{*}_{\mathrm{H}}=\mathop{\arg\min_{\{g_{j}\}\subseteq\mathcal{G}}}R_{\mathrm{H}}^{\ell}(\bm{g})$ denote the minimizer of the corrected risk estimator and the $\ell$-risk w.r.t. the Hamming loss, respectively. Let $\mathfrak{R}_{n,p}(\mathcal{G})$ and $\mathfrak{R}_{n_{j},p_{j}}(\mathcal{G})$ denote the Rademacher complexities defined in Appendix B.4.
Theorem 3.
Assume that the loss function $\ell(z,y)$ is Lipschitz continuous in $z$ with a Lipschitz constant $L_{\ell}$. By the assumptions in Theorem 2, for any $\delta>0$, the following inequality holds with probability at least $1-\delta$:
$R_{\mathrm{H}}^{\ell}(\tilde{\bm{g}}_{\mathrm{H}})-R_{\mathrm{H}}^{\ell}(\bm{g}^{*}_{\mathrm{H}})\leq\frac{8L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}_{n,p}(\mathcal{G})+\frac{16(1-\pi_{j})L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}_{n_{j},p_{j}}(\mathcal{G})$
$+\frac{2}{q}\sum\nolimits_{j=1}^{q}(4-2\pi_{j})C_{\ell}\Delta_{j}+2C_{\ell}\sqrt{\frac{\ln{\left(1/\delta\right)}}{2n}}+\sum\nolimits_{j=1}^{q}\frac{(4-4\pi_{j})C_{\ell}}{q}\sqrt{\frac{\ln{\left(1/\delta\right)}}{2n_{j}}}.$ (11)
The proof can be found in Appendix B.4. Theorem 3 shows that, as $n\rightarrow\infty$, $R_{\mathrm{H}}^{\ell}(\tilde{\bm{g}}_{\mathrm{H}})\rightarrow R_{\mathrm{H}}^{\ell}(\bm{g}^{*}_{\mathrm{H}})$, since $\Delta_{j}\rightarrow 0$, $\mathfrak{R}_{n,p}(\mathcal{G})\rightarrow 0$, and $\mathfrak{R}_{n_{j},p_{j}}(\mathcal{G})\rightarrow 0$ for all parametric models with a bounded norm (Mohri et al., 2012). This means that the minimizer of the corrected risk estimator will approach the desired classifier that minimize the $\ell$-risk w.r.t. the Hamming loss.
Let $R_{\mathrm{H}}^{\ell*}=\mathop{\inf}_{\bm{g}}R_{\mathrm{H}}^{\ell}(\bm{g})$ and $R_{\mathrm{H}}^{*}=\mathop{\inf}_{\bm{f}}R_{\mathrm{H}}^{\mathrm{0-1}}(\bm{f})$ denote the minima of the $\ell$-risk and the risk w.r.t. the Hamming loss, respectively. Then, the following corollary holds.
Corollary 1.
If $\ell$ is a convex function such that $\forall y,\ell^{\prime}(0,y)<0$, then the $\ell$-risk w.r.t. the Hamming loss in Eq. (1) is consistent with the risk w.r.t. the Hamming loss in Eq. (1). This means that, for any sequence of decision functions $\left\{\bm{g}_{t}\right\}$ with corresponding prediction functions $\left\{\bm{f}_{t}\right\}$, if $R^{\ell}_{\mathrm{H}}(\bm{g}_{t})\rightarrow R^{\ell*}_{\mathrm{H}}$, then $R^{\mathrm{0-1}}_{\mathrm{H}}(\bm{f}_{t})\rightarrow R^{*}_{\mathrm{H}}$.
The proof can be found in Appendix B.5. If the model is flexible enough to include the optimal classifier, according to Theorem 3, we have $R^{\ell}_{\mathrm{H}}(\tilde{\bm{g}}_{\mathrm{H}})\rightarrow R^{\ell*}_{\mathrm{H}}$. Then, Corollary 1 demonstrates that $R^{\mathrm{0-1}}_{\mathrm{H}}(\tilde{\bm{f}}_{\mathrm{H}})\rightarrow R^{*}_{\mathrm{H}}$ where $\tilde{\bm{f}}_{\mathrm{H}}$ is the corresponding prediction function of $\tilde{\bm{g}}_{\mathrm{H}}$. This indicates that the prediction function obtained by minimizing the corrected risk estimator in Eq. (3.2) achieves the Bayes risk.
3.3 Second-Order Strategy
The first-order strategy is straightforward but does not consider label correlations, which may be incompatible with multi-label data that exhibit semantic dependencies. Therefore, we explore the ranking loss to model the relationship between pairs of labels. The following theorem applies.
Theorem 4.
When the binary loss function $\ell$ is symmetric, i.e., $\ell(z,\cdot)+\ell(-z,\cdot)=M$ where $M$ is a non-negative constant, then under the assumption in Lemma 1, the $\ell$-risk w.r.t. the ranking loss in Eq. (4) can be equivalently expressed as
$R_{\mathrm{R}}^{\ell}(\bm{g})=$ $\sum\nolimits_{1\leq j<k\leq q}\left((1-\pi_{j})\mathbb{E}_{p(\bm{x}|s_{j}=0)}\left[\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),0\right)\right]\right$.
$\left.+(1-\pi_{k})\mathbb{E}_{p(\bm{x}|s_{k}=0)}\left[\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)\right]-Mp(y_{j}=0,y_{k}=0)\right).$ (12)
The proof can be found in Appendix B.6. Here, the symmetric-loss assumption is often used to ensure statistical consistency of the ranking loss for MLC (Gao and Zhou, 2013). According to Theorem 4, the $\ell$-risk w.r.t. the ranking loss can be expressed as the expectation w.r.t. the conditional density where the $j$-th class label is not regarded as a candidate label. Notably, $Mp(y_{j}=0,y_{k}=0)$ in Eq. (4) is a constant that does not affect training the classifier, so it can be neglected. Similar to the first-order strategy, an unbiased risk estimator can be obtained using $\mathcal{D}_{j}$:
$\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})=$ $\sum\nolimits_{1\leq j<k\leq q}\left(\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}(\bm{x}_{i}^{j})-g_{k}(\bm{x}_{i}^{j}),0\right)\right$.
$\left.+\frac{1-\pi_{k}}{n_{k}}\sum\nolimits_{i=1}^{n_{k}}\ell\left(g_{j}(\bm{x}_{i}^{k})-g_{k}(\bm{x}_{i}^{k}),1\right)\right).$ (13)
To improve generalization performance, we use the flooding regularization technique (Ishida et al., 2020; Liu et al., 2022a; Bae et al., 2024) to mitigate overfitting issues:
$\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})=\left|\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})-\beta\right|+\beta,$ (14)
where $\beta\geq 0$ is a hyper-parameter that controls the minimum of the loss value. Then, we can perform ERM by using Eq. (14). The algorithmic details are summarized in Algorithm 1. We also establish consistency and estimation error bounds for the proposed risk estimator in Eq. (14). The following theorem then holds.
Theorem 5.
Since $R_{\mathrm{R}}^{\ell}(\bm{g})>0$, we assume that there exists a non-negative constant $\gamma$ such that $R_{\mathrm{R}}^{\ell}(\bm{g})\geq\gamma$. We also assume that $\beta$ is chosen such that $\beta\leq\sum\nolimits_{1\leq j<k\leq q}Mp(y_{j}=0,y_{k}=0)$. By the assumptions in Theorem 2, the bias of the expectation of the corrected risk estimator w.r.t. the ranking loss has the following lower and upper bounds:
$0\leq\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\sum\nolimits_{j<k}Mp(y_{j}=0,y_{k}=0)-R_{\mathrm{R}}^{\ell}(\bm{g})\leq\left(2\beta+2C_{\ell}(q-1)\sum\nolimits_{j=1}^{q}(1-\pi_{j})\right)\Delta^{\prime},$ (15)
where $\Delta^{\prime}=\exp{\left(-2\gamma^{2}/\sum_{j=1}^{q}(1-\pi_{j})^{2}(q-1)^{2}C_{\ell}^{2}/n_{j}\right)}$. Furthermore, for any $\delta>0$, the following inequality holds with probability at least $1-\delta$:
$\left|\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})-\sum\nolimits_{j<k}Mp(y_{j}=0,y_{k}=0)-R_{\mathrm{R}}^{\ell}(\bm{g})\right|\leq\sum\nolimits_{j=1}^{q}(1-\pi_{j})(q-1)C_{\ell}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n_{j}}}$
$+\left(2\beta+2C_{\ell}(q-1)\sum\nolimits_{j=1}^{q}(1-\pi_{j})\right)\Delta^{\prime}.$ (16)
The proof can be found in Appendix B.7. According to Theorem 5, as $n\rightarrow\infty$, the bias between the corrected risk estimator in Eq. (14) and the $\ell$-risk of ranking loss will become a constant. This implies that the minimizer of the corrected risk estimator is equivalent to the desired classifier that minimizes the $\ell$-risk w.r.t. the Hamming loss.
Let $\tilde{\bm{g}}_{\mathrm{R}}=\mathop{\arg\min_{\{g_{j}\}\subseteq\mathcal{G}}}\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})$ and $\bm{g}^{*}_{\mathrm{R}}=\mathop{\arg\min_{\{g_{j}\}\subseteq\mathcal{G}}}R_{\mathrm{R}}^{\ell}(\bm{g})$ denote the minimizers of the corrected risk estimator and the $\ell$-risk w.r.t. the ranking loss, respectively.
Theorem 6.
By the assumptions in Theorem 3 and 5, for any $\delta>0$, the following inequality holds with probability at least $1-\delta$:
$R_{\mathrm{R}}^{\ell}(\tilde{\bm{g}}_{\mathrm{R}})-R_{\mathrm{R}}^{\ell}(\bm{g}^{*}_{\mathrm{R}})\leq\left(2\beta+2C_{\ell}(q-1)\sum\nolimits_{j=1}^{q}(1-\pi_{j})\right)\Delta^{\prime}$
$+\sum\nolimits_{j=1}^{q}(1-\pi_{j})(q-1)C_{\ell}\sqrt{\frac{\ln{\left(1/\delta\right)}}{n_{j}}}+\sum\nolimits_{j=1}^{q}4L_{\ell}(q-1)(1-\pi_{j})\mathfrak{R}_{n_{j},p_{j}}\left(\mathcal{G}\right).$ (17)
The proof can be found in Appendix B.8. Theorem 6 shows that as $n\rightarrow\infty$, $R_{\mathrm{R}}^{\ell}(\tilde{\bm{g}}_{\mathrm{R}})\rightarrow R_{\mathrm{R}}^{\ell}(\bm{g}^{*}_{\mathrm{R}})$, since $\Delta^{\prime}\rightarrow 0$ and $\mathfrak{R}_{n_{j},p_{j}}(\mathcal{G})\rightarrow 0$ for all parametric models with a bounded norm (Mohri et al., 2012). This means that the minimizers of Eq. (14) will approach the desired classifiers of the $\ell$-risk w.r.t. the ranking loss when the number of training data increases. Let $R_{\mathrm{R}}^{\ell*}=\mathop{\inf}_{\bm{g}}R_{\mathrm{R}}^{\ell}(\bm{g})$ and $R_{\mathrm{R}}^{*}=\mathop{\inf}_{\bm{f}}R_{\mathrm{R}}^{\mathrm{0-1}}(\bm{f})$ denote the minima of the $\ell$-risk and the risk w.r.t. the ranking loss, respectively. Then we have the following corollary.
Corollary 2.
If $\ell$ is a differentiable, symmetric, and non-increasing function such that $\forall y,\ell^{\prime}(0,y)<0$ and $\ell(z,y)+\ell(-z,y)=M$, then the $\ell$-risk w.r.t. the ranking loss in Eq. (4) is consistent with the risk w.r.t. the ranking loss in Eq. (2.1). This means that for any sequences of decision functions $\left\{\bm{g}_{t}\right\}$ with corresponding prediction functions $\left\{\bm{f}_{t}\right\}$, if $R^{\ell}_{\mathrm{R}}(\bm{g}_{t})\rightarrow R^{\ell*}_{\mathrm{R}}$, then $R^{\mathrm{0-1}}_{\mathrm{R}}(\bm{f}_{t})\rightarrow R^{*}_{\mathrm{R}}$.
The proof can be found in Appendix B.9. If the model is very flexible, we have $R^{\ell}_{\mathrm{R}}(\tilde{\bm{g}}_{\mathrm{R}})\rightarrow R^{\ell*}_{\mathrm{R}}$ according to Theorem 6. Then, Corollary 2 demonstrates that $R^{\mathrm{0-1}}_{\mathrm{R}}(\tilde{\bm{f}}_{\mathrm{R}})\rightarrow R^{*}_{\mathrm{R}}$ where $\tilde{\bm{f}}_{\mathrm{R}}$ is the corresponding prediction function of $\tilde{\bm{g}}_{\mathrm{R}}$. This indicates that the prediction function obtained by minimizing Eq. (14) achieves the Bayes risk.
Ranking Loss $\downarrow$
Approach mirflickr music_emotion music_style yeastBP yeastCC yeastMF
[…]
One Error $\downarrow$
Approach mirflickr music_emotion music_style yeastBP yeastCC yeastMF
[…]
COMES-HL 0.171 $\pm$ 0.019 0.382 $\pm$ 0.015 0.333 $\pm$ 0.012 0.641 $\pm$ 0.030 0.744 $\pm$ 0.020 0.800 $\pm$ 0.023
[…]
Hamming Loss $\downarrow$
Approach mirflickr music_emotion music_style yeastBP yeastCC yeastMF
[…]
Coverage $\downarrow$
Approach mirflickr music_emotion music_style yeastBP yeastCC yeastMF
[…]
Average Precision $\uparrow$
Approach mirflickr music_emotion music_style yeastBP yeastCC yeastMF
[…]
COMES-HL 0.843 $\pm$ 0.013 0.665 $\pm$ 0.009 0.749 $\pm$ 0.010 0.458 $\pm$ 0.020 0.657 $\pm$ 0.020 0.552 $\pm$ 0.023
[…]
Table 2: Experimental results on real-world benchmark datasets. Lower is better for the ranking loss, one error, Hamming loss, coverage; higher is better for the average precision.
4 Experiments
In this section, we validate the effectiveness of the proposed approaches with experimental results.
4.1 Experimental Setup
We conducted experiments on both real-world and synthetic PML benchmark datasets. For real-world datasets, we used mirflickr (Huiskes and Lew, 2008), music_emotion (Zhang and Fang, 2020), music_style, yeastBP (Yu et al., 2018a), yeastCC and yeastMF; for synthetic datasets, we used VOC2007 (Everingham et al., 2007), VOC2012 (Everingham et al., 2012), CUB (Wah et al., 2011) and COCO2014 (Lin et al., 2014), where candidate labels were generated by two different data generation processes. Full experimental details are given in Appendix C. Following standard practice (Liu et al., 2023), we evaluated with ranking loss, one error, Hamming loss, coverage and average precision on real-world sets, and with mean average precision (mAP) on synthetic datasets.
4.2 Experimental Results
Tables 2 and 3 summarize results on real-world and synthetic datasets, respectively. Here $\bullet$ indicates that the best method is significantly better than its competitor (paired $t$-test at 0.05 significance level). We observe that both instantiations of COMES consistently outperform other baselines across various datasets, clearly validating the effectiveness of our proposed approaches. We attribute this to: (1) our data generation assumptions are more realistic and better match statistics of real-world datasets; (2) our risk-correction approaches effectively mitigates overfitting, an issue overlooked by previous unbiased methods (Xie and Huang, 2023).
4.3 Sensitivity Analysis
Influence of Inaccurate Class Priors. To investigate the influence of inaccurate class priors, we added Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ to each class prior $\pi_{j}$. The experimental results are shown in Figure 2. We observe that COMES-HL is more sensitive to inaccurate class priors. Overall performance remains stable within a reasonable range of class priors, but may degrade when the priors become highly inaccurate.
VOC2007 VOC2012 CUB COCO2014
Approach Case-a Case-b Case-a Case-b Case-a Case-b Case-a Case-b
[…]
COMES-HL 42.33 $\pm$ 1.74 42.43 $\pm$ 4.17 48.72 $\pm$ 1.08 47.93 $\pm$ 1.05 18.94 $\pm$ 0.30 18.95 $\pm$ 0.39 33.62 $\pm$ 0.57 32.76 $\pm$ 1.45
[…]
Table 3: Classification performance in terms of mAP on synthetic benchmark datasets.
Figure 2: Classification performance with inaccurate class priors on different datasets.
Figure 3: Sensitivity analysis w.r.t. $\beta$ on the mirflickr dataset.
Influence of $\beta$. We also investigated the influence of the hyperparameter $\beta$ for the flooding regularization used in COMES-RL. From Figure 3, we observe that the performance of COMES-RL is rather stable when $\beta$ is set within a reasonable range. During our experiments, we found that the performance was already competitive by setting $\beta=0$ on many datasets. However, the performance may degrade when $\beta$ is set to a large value. This also matches our theoretical results in Theorem 5, where consistency holds when $\beta$ is not too large.
5 Conclusion
In this paper, we rethought MLC under inexact supervision by proposing a novel framework. We proposed two instantiations of unbiased risk estimators w.r.t. the Hamming loss and ranking loss, two widely used evaluation metrics for MLC, respectively. We also introduced risk-correction approaches to improve generalization performance with theoretical guarantees. Extensive experiments on ten real-world and synthetic benchmark datasets validated the effectiveness of the proposed approaches. A limitation of this work is that we consider the generation process to be independent of the instances. In the future, it is promising to extend our proposed methodologies to instance-dependent settings.
Appendix A More Details about the Algorithm
A.1 Algorithmic Pseudo-Code
Algorithm 1 COMES-HL and COMES-RL
Input: Multi-label classifiers $\bm{g}$, PML dataset $\mathcal{D}$, epoch number $T_{\text{max}}$, iteration number $I_{\text{max}}$.
1:for $t$ = $1,2,\ldots,T_{\text{max}}$ do
2:  Shuffle $\mathcal{D}$;
3:  for $j=1,\ldots,I_{\text{max}}$ do
4:   Fetch mini-batch $\mathcal{D}_{j}$ from $\mathcal{D}$;
5:   Forward $\mathcal{D}$ and get the outputs of $\bm{g}$;
6:   if using the COMES-HL algorithm then
7:     Calculating the loss based on Eq. (3.2);
8:   else if using the COMES-HL algorithm then
9:     Calculating the loss based on Eq. (14);
10:   end if
11:   Update $\bm{g}$ using a stochastic optimizer to minimize the loss;
12:  end for
13:end for
Output: $\bm{g}$.
A.2 Class Prior Estimation
We can use any off-the-shelf mixture proportion estimation algorithm to estimate the class priors [Ramaswamy et al., 2016, Garg et al., 2021, Yao et al., 2022, Zhu et al., 2023], which are mainly designed to estimate the class prior with positive and unlabeled data for binary classification. Specifically, we generate negative and unlabeled datasets according to Eq. (6) and then apply the mixture proportion estimation algorithm. The algorithmic details are summarized in Algorithm 2.
Algorithm 2 Class-prior Estimation
Input: Mixture proportion estimation algorithm $\mathcal{A}$, PML dataset $\mathcal{D}$.
1:for $k\in\mathcal{Y}$ do
2:  Generate unlabeled and negative datasets according to Eq. (6);
3:  Estimate the value of $(1-\pi_{k})$ by using $\mathcal{A}$ and interchanging the positive and negative classes;
4:end for
Output: Class priors $\pi_{k}$ ($k\in\mathcal{Y}$).
Appendix B Proofs
B.1 Proof of Lemma 1
On one hand, we have
$p(\bm{x}|s_{j}=0,j\notin Y)=\frac{p\left(\bm{x}|s_{j}=0\right)p(j\notin Y|\bm{x},s_{j}=0)}{p(j\notin Y|s_{j}=0)}=p\left(\bm{x}|s_{j}=0\right).$
On the other hand, we have
$p(\bm{x}|s_{j}=0,j\notin Y)=\frac{p(\bm{x}|j\notin Y)p(s_{j}=0|\bm{x},j\notin Y)}{p(s_{j}=0|j\notin Y)}=p(\bm{x}|j\notin Y),$
where the second equation is due to $p(s_{j}=0|\bm{x},j\notin Y)=p(s_{j}=0|j\notin Y)=p_{j}$. The proof is completed. ∎
B.2 Proof of Theorem 1
$R_{\mathrm{H}}^{\ell}(\bm{g})=$ $\mathbb{E}_{p(\bm{x},Y)}\left[\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),y_{j}\right)\right]$
$=$ $\int\sum\nolimits_{Y}\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),y_{j}\right)p\left(\bm{x},Y\right)\mathop{}\!\mathrm{d}\bm{x}$
[…]
$=$ $\int\frac{1}{q}\sum\nolimits_{j=1}^{q}\sum\nolimits_{y_{j}=0}^{1}\ell\left(g_{j}\left(\bm{x}\right),y_{j}\right)p(\bm{x},y_{j})\mathop{}\!\mathrm{d}\bm{x}$
$=$ $\int\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),1\right)p(\bm{x},y_{j}=1)\mathop{}\!\mathrm{d}\bm{x}+\int\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),0\right)p\left(\bm{x},y_{j}=0\right)\mathop{}\!\mathrm{d}\bm{x}$
$=$ $\int\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),1\right)p(\bm{x})\mathop{}\!\mathrm{d}\bm{x}-\int\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),1\right)p(\bm{x},y_{j}=0)\mathop{}\!\mathrm{d}\bm{x}$
$+\int\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),0\right)p\left(\bm{x},y_{j}=0\right)\mathop{}\!\mathrm{d}\bm{x}$
$=$ $\int\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),1\right)p(\bm{x})\mathop{}\!\mathrm{d}\bm{x}+\int\frac{1}{q}\sum\nolimits_{j=1}^{q}\left(\ell\left(g_{j}\left(\bm{x}\right),0\right)-\ell\left(g_{j}\left(\bm{x}\right),1\right)\right)\left(1-\pi_{j}\right)p\left(\bm{x}|y_{j}=0\right)\mathop{}\!\mathrm{d}\bm{x}$
$=$ $\mathbb{E}_{p(\bm{x})}\left[\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),1\right)\right]+\mathbb{E}_{p(\bm{x}|y_{j}=0)}\left[\frac{1}{q}\sum\nolimits_{j=1}^{q}\left(1-\pi_{j}\right)\left(\ell\left(g_{j}\left(\bm{x}\right),0\right)-\ell\left(g_{j}\left(\bm{x}\right),1\right)\right)\right]$
$=$ $\mathbb{E}_{p(\bm{x})}\left[\frac{1}{q}\sum\nolimits_{j=1}^{q}\ell\left(g_{j}\left(\bm{x}\right),1\right)\right]+\mathbb{E}_{p(\bm{x}|s_{j}=0)}\left[\frac{1}{q}\sum\nolimits_{j=1}^{q}\left(1-\pi_{j}\right)\left(\ell\left(g_{j}\left(\bm{x}\right),0\right)-\ell\left(g_{j}\left(\bm{x}\right),1\right)\right)\right],$
where the last equation is by Lemma 1. The proof is completed. ∎
B.3 Proof of Theorem 2
Let
$\mathfrak{D}^{+}_{j}(g_{j})=\left\{\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)|\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)>0\right\}$
and
$\mathfrak{D}^{-}_{j}(g_{j})=\left\{\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)|\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\leq 0\right\}$
denote the set of data pairs with positive and negative empirical losses, respectively. Then, we have the following lemma.
Lemma 2.
The probability measure of $\mathfrak{D}^{-}_{j}(g_{j})$ can be bounded as follows:
$\mathbb{P}\left(\mathfrak{D}^{-}_{j}(g_{j})\right)\leq\exp\left(\frac{-2\alpha^{2}}{C_{\ell}^{2}/n+(1-\pi_{j})^{2}C_{\ell}^{2}/n_{j}}\right).$ (18)
Proof.
Let
$p\left(\mathcal{D}_{\rm U}\right)=p(\bm{x}_{1}^{\rm U})p(\bm{x}_{2}^{\rm U})\cdots p(\bm{x}_{n}^{\rm U})\quad\text{and}\quad p\left(\mathcal{D}_{j}\right)=p(\bm{x}_{1}^{j})p(\bm{x}_{2}^{j})\cdots p(\bm{x}_{n_{j}}^{j})$
denote the densities of $\mathcal{D}_{\rm U}$ and $\mathcal{D}_{j}$, respectively. Then, the joint density of $\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)$ is
$p\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)=p\left(\mathcal{D}_{\rm U}\right)p\left(\mathcal{D}_{j}\right).$
Then, the probability measure of $\mathfrak{D}^{-}_{j}(g_{j})$ can be expressed as
$\mathbb{P}\left(\mathfrak{D}^{-}_{j}(g_{j})\right)=$ $\int_{\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)\in\mathfrak{D}^{-}_{j}(g_{j})}p\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)\mathop{}\!\mathrm{d}\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)$
$=$ $\int_{\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)\in\mathfrak{D}^{-}_{j}(g_{j})}p\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)\mathop{}\!\mathrm{d}\bm{x}_{1}^{\rm U}\mathop{}\!\mathrm{d}\bm{x}_{2}^{\rm U}\cdots\mathop{}\!\mathrm{d}\bm{x}_{n}^{\rm U}\mathop{}\!\mathrm{d}\bm{x}_{1}^{j}\mathop{}\!\mathrm{d}\bm{x}_{2}^{j}\cdots\mathop{}\!\mathrm{d}\bm{x}_{n_{j}}^{j}$
When an instance in $\mathcal{D}_{\rm U}$ is replaced by another instance, the value of $\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)/n-(1-\pi_{j})\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)/n_{j}$ changes no more than $C_{\ell}/n$. When an instance in $\mathcal{D}_{j}$ is replaced by another instance, the value of $\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)/n-(1-\pi_{j})\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)/n_{j}$ changes no more than $(1-\pi_{j})C_{\ell}/n_{j}$. Therefore, by applying the McDiarmid’s inequality, we can obtain the following inequality:
$p\left(\pi_{j}\mathbb{E}_{p(\bm{x}|y_{j}=1)}\left[\ell\left(g_{j}\left(\bm{x}\right),1\right)\right]-\left(\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\right)\geq\alpha\right)$
$\leq\exp\left(\frac{-2\alpha^{2}}{C_{\ell}^{2}/n+(1-\pi_{j})^{2}C_{\ell}^{2}/n_{j}}\right).$
Then we have
$\mathbb{P}\left(\mathfrak{D}^{-}_{j}(g_{j})\right)=$ $p\left(\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\leq 0\right)$
$\leq$ $p\left(\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\leq\pi_{j}\mathbb{E}_{p(\bm{x}|y_{j}=1)}\left[\ell\left(g_{j}\left(\bm{x}\right),1\right)\right]-\alpha\right)$
$\leq$ $\exp\left(\frac{-2\alpha^{2}}{C_{\ell}^{2}/n+(1-\pi_{j})^{2}C_{\ell}^{2}/n_{j}}\right),$
which concludes the proof. ∎
Then, we provide the proof of Theorem 2.
Proof of Theorem 2.
To begin with, we have
$\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-R_{\mathrm{H}}^{\ell}(\bm{g})=\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})-\hat{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]\geq 0.$
Besides, we have
$\left|\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\right|$
$\leq$ $\left|\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)\right|+\left|\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\right|$
$\leq$ $(2-\pi_{j})C_{\ell}.$
Then,
$\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-R_{\mathrm{H}}^{\ell}(\bm{g})$
$=$ $\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})-\hat{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]$
$=$ $\frac{1}{q}\sum\nolimits_{j=1}^{q}\int_{\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)\in\mathfrak{D}^{-}_{j}(g_{j})}\left(\left|\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\right|\right$.
$\left.-\left(\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\right)\right)p\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)\mathop{}\!\mathrm{d}\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)$
$\leq$ $\frac{1}{q}\sum\nolimits_{j=1}^{q}\sup_{\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)\in\mathfrak{D}^{-}_{j}(g_{j})}\left(\left|\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\right|\right$.
$\left.-\left(\frac{1}{n}\sum\nolimits_{i=1}^{n}\ell\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\ell\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\right)\right)\int_{\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)\in\mathfrak{D}^{-}_{j}(g_{j})}p\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)\mathop{}\!\mathrm{d}\left(\mathcal{D}_{\rm U},\mathcal{D}_{j}\right)$
$\leq$ $\frac{1}{q}\sum\nolimits_{j=1}^{q}(4-2\pi_{j})C_{\ell}\mathbb{P}\left(\mathfrak{D}^{-}_{j}(g_{j})\right)$
$\leq$ $\frac{1}{q}\sum\nolimits_{j=1}^{q}(4-2\pi_{j})C_{\ell}\exp\left(\frac{-2\alpha^{2}}{C_{\ell}^{2}/n+(1-\pi_{j})^{2}C_{\ell}^{2}/n_{j}}\right)$
$=$ $\frac{1}{q}\sum\nolimits_{j=1}^{q}(4-2\pi_{j})C_{\ell}\Delta_{j},$
which concludes the proof of the first part of the theorem. Then, we provide an upper bound for $\left|\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|$. When an instance in $\mathcal{D}_{\rm U}$ is replaced by another instance, the value of $\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})$ changes at most $C_{\ell}/n$; when an instance in $\mathcal{D}_{j}$ is replaced by another instance, the value of $\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})$ changes at most $(2-2\pi_{j})C_{\ell}/(qn_{j})$. By applying McDiarmid’s inequality, we have the following inequalities with probability at least $1-\delta/2$:
$\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\leq C_{\ell}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n}}+\sum\nolimits_{j=1}^{q}\frac{(2-2\pi_{j})C_{\ell}}{q}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n_{j}}},$
$\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})-\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]\leq C_{\ell}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n}}+\sum\nolimits_{j=1}^{q}\frac{(2-2\pi_{j})C_{\ell}}{q}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n_{j}}},$
where we use the inequality that $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$. Therefore, the following inequality holds with probability at least $1-\delta$:
$\left|\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|\leq C_{\ell}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n}}+\sum\nolimits_{j=1}^{q}\frac{(2-2\pi_{j})C_{\ell}}{q}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n_{j}}}.$
Finally, we have
$\left|\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})-R_{\mathrm{H}}^{\ell}(\bm{g})\right|$
$=$ $\left|\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})-\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]+\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-R_{\mathrm{H}}^{\ell}(\bm{g})\right|$
$\leq$ $\left|\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})-\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]\right|+\left|\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-R_{\mathrm{H}}^{\ell}(\bm{g})\right|$
$=$ $\left|\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})-\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]\right|+\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-R_{\mathrm{H}}^{\ell}(\bm{g})$
$\leq$ $\frac{1}{q}\sum\nolimits_{j=1}^{q}\left((4-2\pi_{j})C_{\ell}\exp\left(\frac{-2\alpha^{2}}{C_{\ell}^{2}/n+(1-\pi_{j})^{2}C_{\ell}^{2}/n_{j}}\right)+\frac{(2-2\pi_{j})C_{\ell}}{q}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n_{j}}}\right)+C_{\ell}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n}}$
$=$ $\frac{1}{q}\sum\nolimits_{j=1}^{q}\left((4-2\pi_{j})C_{\ell}\Delta_{j}+\frac{(2-2\pi_{j})C_{\ell}}{q}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n_{j}}}\right)+C_{\ell}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n}},$
which concludes the proof. ∎
B.4 Proof of Theorem 3
Definition 1 (Rademacher complexity).
Let $\mathcal{X}_{n}=\{\bm{x}_{1},\ldots\bm{x}_{n}\}$ denote $n$ i.i.d. random variables drawn from a probability distribution with density $p(\bm{x})$, $\mathcal{G}=\{g_{k}:\mathcal{X}\mapsto\mathbb{R}\}$ denote a class of measurable functions of model outputs for the $k$,-th class, and $\bm{\sigma}=(\sigma_{1},\sigma_{2},\ldots,\sigma_{n})$ denote Rademacher variables taking values from $\{+1,-1\}$ uniformly. Then, the (expected) Rademacher complexity of $\mathcal{G}$ is defined as
$\mathfrak{R}_{n,p}(\mathcal{G})=\mathbb{E}_{\mathcal{X}_{n}}\mathbb{E}_{\bm{\sigma}}\left[\sup_{g_{j}\in\mathcal{G}}\frac{1}{n}\sum\nolimits_{i=1}^{n}\sigma_{i}g_{j}(\bm{x}_{i})\right].$ (19)
We also introduce an alternative definition of Rademacher complexity:
$\mathfrak{R}^{\prime}_{n,p}(\mathcal{G})=\mathbb{E}_{\mathcal{X}_{n}}\mathbb{E}_{\bm{\sigma}}\left[\sup_{g_{j}\in\mathcal{G}}\left|\frac{1}{n}\sum\nolimits_{i=1}^{n}\sigma_{i}g_{j}(\bm{x}_{i})\right|\right].$ (20)
Then, we introduce the following lemmas.
Lemma 3.
Without any composition, for any $\mathcal{G}$, we have $\mathfrak{R}^{\prime}_{n,p}(\mathcal{G})\geq\mathfrak{R}_{n,p}(\mathcal{G})$. If $\mathcal{G}$ is closed under negation, we have $\mathfrak{R}^{\prime}_{n,p}(\mathcal{G})=\mathfrak{R}_{n,p}(\mathcal{G})$.
Lemma 4 (Theorem 4.12 in [Ledoux and Talagrand, 1991]).
If $\ell:\mathbb{R}\times\left\{0,1\right\}\rightarrow\mathbb{R}$ is a Lipschitz continuous function with a Lipschitz constant $L_{\ell}$ and satisfies $\forall y,\ell(0,y)=0$, we have
$\mathfrak{R}^{\prime}_{n,p}(\ell\circ\mathcal{G})\leq 2L_{\ell}\mathfrak{R}^{\prime}_{n,p}(\mathcal{G}),$
where $\ell\circ\mathcal{G}=\{\ell\circ g_{j}|g_{j}\in\mathcal{G}\}$.
Then, we provide the following lemma.
Lemma 5.
Based on the above assumptions, for any $\delta>0$, the following inequality holds with probability at least $1-\delta$:
$\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|R_{\mathrm{H}}^{\ell}(\bm{g})-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|\leq\frac{4L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}^{\prime}_{n,p}(\mathcal{G})+\frac{8(1-\pi_{j})L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}^{\prime}_{n_{j},p_{j}}(\mathcal{G})$
$+\frac{1}{q}\sum\nolimits_{j=1}^{q}(4-2\pi_{j})C_{\ell}\Delta_{j}+C_{\ell}\sqrt{\frac{\ln{\left(1/\delta\right)}}{2n}}+\sum\nolimits_{j=1}^{q}\frac{(2-2\pi_{j})C_{\ell}}{q}\sqrt{\frac{\ln{\left(1/\delta\right)}}{2n_{j}}}.$
Proof.
When an instance in $\mathcal{D}_{\rm U}$ is replaced by another instance, the value of $\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|$ changes at most $C_{\ell}/n$; when an instance in $\mathcal{D}_{j}$ is replaced by another instance, the value of $\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|$ changes at most $(2-2\pi_{j})C_{\ell}/(qn_{j})$. By applying McDiarmid’s inequality, we have the following inequality with probability at least $1-\delta$:
$\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|-\mathbb{E}\left[\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|\right]$
$\leq$ $C_{\ell}\sqrt{\frac{\ln{\left(1/\delta\right)}}{2n}}+\sum\nolimits_{j=1}^{q}\frac{(2-2\pi_{j})C_{\ell}}{q}\sqrt{\frac{\ln{\left(1/\delta\right)}}{2n_{j}}}.$ (21)
For ease of notations, let $\overline{\mathcal{D}}=\mathcal{D}_{\rm U}\bigcup\mathcal{D}_{1}\bigcup\mathcal{D}_{2}\ldots\bigcup\mathcal{D}_{q}$ denote set of all the data. We have
$\mathbb{E}\left[\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|\right]$
$=$ $\mathbb{E}_{\overline{\mathcal{D}}}\left[\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}_{\overline{\mathcal{D}}^{\prime}}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|\right]$
$\leq$ $\mathbb{E}_{\overline{\mathcal{D}},\overline{\mathcal{D}}^{\prime}}\left[\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g};\overline{\mathcal{D}})-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g}^{\prime};\overline{\mathcal{D}}^{\prime})\right|\right],$ (22)
[…]
$\left|\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g};\widehat{\mathcal{D}})-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g};\widehat{\mathcal{D}}^{\prime})\right|$
[…]
$\left.\left|\frac{1}{n}\sum\nolimits_{i=1}^{n}\bar{\ell}\left(g_{j}\left(\bm{x}^{\mathrm{U}^{\prime}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\bar{\ell}\left(g_{j}\left(\bm{x}^{j^{\prime}}_{i}\right),1\right)\right|\right|$
$+\sum\nolimits_{j=1}^{q}\left|\frac{1-\pi_{j}}{qn_{j}}\sum\nolimits_{i=1}^{n_{j}}\bar{\ell}\left(g_{j}\left(\bm{x}^{j}_{i}\right),0\right)-\frac{1-\pi_{j}}{qn_{j}}\sum\nolimits_{i=1}^{n_{j}}\bar{\ell}\left(g_{j}\left(\bm{x}^{j^{\prime}}_{i}\right),0\right)\right|$
$\leq$ $\frac{1}{q}\sum\nolimits_{j=1}^{q}\left|\frac{1}{n}\sum\nolimits_{i=1}^{n}\bar{\ell}\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\bar{\ell}\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\right$.
$\left.-\frac{1}{n}\sum\nolimits_{i=1}^{n}\bar{\ell}\left(g_{j}\left(\bm{x}^{\mathrm{U}^{\prime}}_{i}\right),1\right)+\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\bar{\ell}\left(g_{j}\left(\bm{x}^{j^{\prime}}_{i}\right),1\right)\right|$
$+\sum\nolimits_{j=1}^{q}\left|\frac{1-\pi_{j}}{qn_{j}}\sum\nolimits_{i=1}^{n_{j}}\bar{\ell}\left(g_{j}\left(\bm{x}^{j}_{i}\right),0\right)-\frac{1-\pi_{j}}{qn_{j}}\sum\nolimits_{i=1}^{n_{j}}\bar{\ell}\left(g_{j}\left(\bm{x}^{j^{\prime}}_{i}\right),0\right)\right|$
$\leq$ $\frac{1}{q}\sum\nolimits_{j=1}^{q}\left|\frac{1}{n}\sum\nolimits_{i=1}^{n}\bar{\ell}\left(g_{j}\left(\bm{x}^{\mathrm{U}}_{i}\right),1\right)-\frac{1}{n}\sum\nolimits_{i=1}^{n}\bar{\ell}\left(g_{j}\left(\bm{x}^{\mathrm{U}^{\prime}}_{i}\right),1\right)\right|$
$+\frac{1}{q}\sum\nolimits_{j=1}^{q}\left|\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\bar{\ell}\left(g_{j}\left(\bm{x}^{j^{\prime}}_{i}\right),1\right)-\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\bar{\ell}\left(g_{j}\left(\bm{x}^{j}_{i}\right),1\right)\right|$
$+\sum\nolimits_{j=1}^{q}\left|\frac{1-\pi_{j}}{qn_{j}}\sum\nolimits_{i=1}^{n_{j}}\bar{\ell}\left(g_{j}\left(\bm{x}^{j}_{i}\right),0\right)-\frac{1-\pi_{j}}{qn_{j}}\sum\nolimits_{i=1}^{n_{j}}\bar{\ell}\left(g_{j}\left(\bm{x}^{j^{\prime}}_{i}\right),0\right)\right|,$ (23)
where the inequalities are due to the triangle inequality. Then, by combining Inequalities (22) and (23), it is a routine work [Mohri et al., 2012] to show that
$\mathbb{E}_{\overline{\mathcal{D}},\overline{\mathcal{D}}^{\prime}}\left[\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g};\overline{\mathcal{D}})-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g}^{\prime};\overline{\mathcal{D}}^{\prime})\right|\right]$
$\leq$ $\frac{2}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}^{\prime}_{n,p}(\bar{\ell}\circ\mathcal{G})+\frac{4(1-\pi_{j})}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}^{\prime}_{n_{j},p_{j}}(\bar{\ell}\circ\mathcal{G})$
$\leq$ $\frac{4L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}^{\prime}_{n,p}(\mathcal{G})+\frac{8(1-\pi_{j})L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}^{\prime}_{n_{j},p_{j}}(\mathcal{G})$
$=$ $\frac{4L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}_{n,p}(\mathcal{G})+\frac{8(1-\pi_{j})L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}_{n_{j},p_{j}}(\mathcal{G}),$ (24)
where the second inequality is due to Lemma 4, $p_{j}$ denotes $p(\bm{x}|s_{j}=0)$, and the last equality is due to Lemma 3. By combining Inequalities (B.4) and (B.4), we have the following inequality with probability at least $1-\delta$:
$\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|\leq\frac{4L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}_{n,p}(\mathcal{G})+\frac{8(1-\pi_{j})L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}_{n_{j},p_{j}}(\mathcal{G})$
$+C_{\ell}\sqrt{\frac{\ln{\left(1/\delta\right)}}{2n}}+\sum\nolimits_{j=1}^{q}\frac{(2-2\pi_{j})C_{\ell}}{q}\sqrt{\frac{\ln{\left(1/\delta\right)}}{2n_{j}}}.$ (25)
Then, we have the following inequality with probability at least $1-\delta$:
$\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|R_{\mathrm{H}}^{\ell}(\bm{g})-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|$
$=$ $\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|R_{\mathrm{H}}^{\ell}(\bm{g})-\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]+\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|$
$\leq$ $\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|R_{\mathrm{H}}^{\ell}(\bm{g})-\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]\right|+\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|$
$\leq$ $\frac{1}{q}\sum\nolimits_{j=1}^{q}(4-2\pi_{j})C_{\ell}\Delta_{j}+C_{\ell}\sqrt{\frac{\ln{\left(1/\delta\right)}}{2n}}+\sum\nolimits_{j=1}^{q}\frac{(2-2\pi_{j})C_{\ell}}{q}\sqrt{\frac{\ln{\left(1/\delta\right)}}{2n_{j}}}$
$+\frac{4L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}_{n,p}(\mathcal{G})+\frac{8(1-\pi_{j})L_{\ell}}{q}\sum\nolimits_{j=1}^{q}\mathfrak{R}_{n_{j},p_{j}}(\mathcal{G}),$
where the second inequality is due to Inequalities B.4 and 9. The proof is complete. ∎
Then, we provide the proof of Theorem 3.
Proof of Theorem 3.
$R_{\mathrm{H}}^{\ell}(\tilde{\bm{g}}_{\mathrm{H}})-R_{\mathrm{H}}^{\ell}(\bm{g}^{*}_{\mathrm{H}})$ $=R_{\mathrm{H}}^{\ell}(\tilde{\bm{g}}_{\mathrm{H}})-\tilde{R}_{\mathrm{H}}^{\ell}(\tilde{\bm{g}}_{\mathrm{H}})+\tilde{R}_{\mathrm{H}}^{\ell}(\tilde{\bm{g}}_{\mathrm{H}})-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g}^{*}_{\mathrm{H}})+\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g}^{*}_{\mathrm{H}})-R_{\mathrm{H}}^{\ell}(\bm{g}^{*}_{\mathrm{H}})$
$\leq R_{\mathrm{H}}^{\ell}(\tilde{\bm{g}}_{\mathrm{H}})-\tilde{R}_{\mathrm{H}}^{\ell}(\tilde{\bm{g}}_{\mathrm{H}})+\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g}^{*}_{\mathrm{H}})-R_{\mathrm{H}}^{\ell}(\bm{g}^{*}_{\mathrm{H}})$
$\leq 2\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|R_{\mathrm{H}}^{\ell}(\bm{g})-\tilde{R}_{\mathrm{H}}^{\ell}(\bm{g})\right|.$
By Lemma 5, the proof is complete. ∎
B.5 Proof of Corollary 1
Lemma 6 (Theorem 4 in Gao and Zhou [2013]).
The surrogate loss $R^{\ell}$ is multi-label consistent w.r.t. the Hamming or ranking loss $R^{\mathrm{0-1}}$ if and only if it holds for any sequence $\left\{\bm{g}_{t}\right\}$ that if $R^{\ell}(\bm{g})\rightarrow R^{\ell*}$, then $R^{\mathrm{0-1}}(\bm{g})\rightarrow R^{*}$. Here, $R^{\ell*}=\mathop{\inf}_{\bm{g}}R^{\ell}(\bm{g})$ and $R^{*}=\mathop{\inf}_{\bm{g}}R^{\mathrm{0-1}}(\bm{g})$.
Lemma 7 (Theorem 32 in Gao and Zhou [2013]).
If $\ell$ is a convex function with $\ell^{\prime}(0,y)<0$, then Eq. (2) is consistent w.r.t. the Hamming loss.
Then, we provide the proof of Corollary 1.
Proof of Corollary 1.
Since the proposed risk in Eq. (1) is equivalent to the risk in Eq. (2), it is sufficient to prove that for any sequence $\left\{\bm{g}_{t}\right\}$ that if $R^{\ell}_{\mathrm{H}}(\bm{g}_{t})\rightarrow R^{\ell*}_{\mathrm{H}}$, then $R^{\mathrm{0-1}}_{\mathrm{H}}(\bm{f}_{t})\rightarrow R^{*}_{\mathrm{H}}$. ∎
B.6 Proof of Theorem 4
$R_{\mathrm{R}}^{\ell}(\bm{g})=$ $\mathbb{E}_{p(\bm{x},Y)}\left[\sum\nolimits_{1\leq j<k\leq q}\mathbb{I}(y_{j}\neq y_{k})\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),\frac{y_{j}-y_{k}+1}{2}\right)\right]$
$=$ $\int\sum\nolimits_{Y}\sum\nolimits_{1\leq j<k\leq q}\mathbb{I}(y_{j}\neq y_{k})\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),\frac{y_{j}-y_{k}+1}{2}\right)p\left(\bm{x},Y\right)\mathop{}\!\mathrm{d}\bm{x}$
$=$ $\int\sum\nolimits_{1\leq j<k\leq q}\sum\nolimits_{y_{j}=0}^{1}\sum\nolimits_{y_{k}=0}^{1}\sum\nolimits_{Y^{\prime}=Y\backslash\{y_{j},y_{k}\}}\mathbb{I}(y_{j}\neq y_{k})\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),\frac{y_{j}-y_{k}+1}{2}\right)$
$p\left(\bm{x},y_{j},y_{k}\right)p\left(Y^{\prime}|\bm{x},y_{j},y_{k}\right)\mathop{}\!\mathrm{d}\bm{x}$
$=$ $\int\sum\nolimits_{1\leq j<k\leq q}\sum\nolimits_{y_{j}=0}^{1}\sum\nolimits_{y_{k}=0}^{1}\mathbb{I}(y_{j}\neq y_{k})\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),\frac{y_{j}-y_{k}+1}{2}\right)$
$p\left(\bm{x},y_{j},y_{k}\right)\sum\nolimits_{Y^{\prime}=Y\backslash\{y_{j},y_{k}\}}p\left(Y^{\prime}|\bm{x},y_{j},y_{k}\right)\mathop{}\!\mathrm{d}\bm{x}$
$=$ $\int\sum\nolimits_{1\leq j<k\leq q}\sum\nolimits_{y_{j}=0}^{1}\sum\nolimits_{y_{k}=0}^{1}\mathbb{I}(y_{j}\neq y_{k})\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),\frac{y_{j}-y_{k}+1}{2}\right)p\left(\bm{x},y_{j},y_{k}\right)\mathop{}\!\mathrm{d}\bm{x}$
$=$ $\sum\nolimits_{1\leq j<k\leq q}\left(\int\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),0\right)p\left(\bm{x},y_{j}=0,y_{k}=1\right)\mathop{}\!\mathrm{d}\bm{x}\right$.
$\left.+\int\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)p\left(\bm{x},y_{j}=1,y_{k}=0\right)\mathop{}\!\mathrm{d}\bm{x}\right)$
$=$ $\sum\nolimits_{1\leq j<k\leq q}\left(\int\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),0\right)\left(p\left(\bm{x},y_{j}=0\right)-p\left(\bm{x},y_{j}=0,y_{k}=0\right)\right)\mathop{}\!\mathrm{d}\bm{x}\right$.
$\left.+\int\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)\left(p\left(\bm{x},y_{k}=0\right)-p\left(\bm{x},y_{j}=0,y_{k}=0\right)\right)\mathop{}\!\mathrm{d}\bm{x}\right)$
$=$ $\sum\nolimits_{1\leq j<k\leq q}\left(\int\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),0\right)(1-\pi_{j})p\left(\bm{x}|y_{j}=0\right)\mathop{}\!\mathrm{d}\bm{x}\right$.
$+\int\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)(1-\pi_{k})p\left(\bm{x}|y_{k}=0\right)\mathop{}\!\mathrm{d}\bm{x}$
$\left.-\int\left(\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)+\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)\right)p\left(\bm{x},y_{j}=0,y_{k}=0\right)\mathop{}\!\mathrm{d}\bm{x}\right)$
$=$ $\sum\nolimits_{1\leq j<k\leq q}\left((1-\pi_{j})\mathbb{E}_{p(\bm{x}|y_{j}=0)}\left[\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),0\right)\right]\right$.
$\left.+(1-\pi_{k})\mathbb{E}_{p(\bm{x}|y_{k}=0)}\left[\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)\right]-Cp(y_{j}=0,y_{k}=0)\right).$
$=$ $\sum\nolimits_{1\leq j<k\leq q}\left((1-\pi_{j})\mathbb{E}_{p(\bm{x}|s_{j}=0)}\left[\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),0\right)\right]\right$.
$\left.+(1-\pi_{k})\mathbb{E}_{p(\bm{x}|s_{k}=0)}\left[\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)\right]-Cp(y_{j}=0,y_{k}=0)\right),$
where the last equation is by Lemma 1. The proof is completed. ∎
B.7 Proof of Theorem 5
Let $\widehat{\mathcal{D}}=\mathcal{D}_{1}\bigcup\mathcal{D}_{2}\bigcup\ldots\mathcal{D}_{q}$ denote the set of all the data used in Eq. (14). We introduce
$\mathfrak{D}^{+}(\bm{g})=\left\{\widehat{\mathcal{D}}|\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})>\beta\right\},\quad\mathrm{and}\quad\mathfrak{D}^{-}(\bm{g})=\left\{\widehat{\mathcal{D}}|\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})\leq\beta\right\}.$
Then, we have the following lemma.
Lemma 8.
The probability measure of $\mathfrak{D}^{-}(\bm{g})$ can be bounded as follows:
$\mathbb{P}\left(\mathfrak{D}^{-}(\bm{g})\right)\leq\exp\left(\frac{-2\gamma^{2}}{\sum_{j=1}^{q}(1-\pi_{j})^{2}(q-1)^{2}C_{\ell}^{2}/n_{j}}\right).$ (26)
Proof.
When an instance from $\mathcal{D}_{j}$ is replaced by another instance, the value of $\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})$ changes at most $(1-\pi_{j})(q-1)C_{\ell}/n_{j}$. Therefore, by applying the McDiarmid’s inequality, we can obtain the following inequality:
$p\left(R_{\mathrm{R}}^{\ell}(\bm{g})-\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})+\sum\nolimits_{1\leq j<k\leq q}Mp(y_{j}=0,y_{k}=0)\geq\gamma\right)\leq\exp\left(\frac{-2\gamma^{2}}{\sum_{j=1}^{q}(1-\pi_{j})^{2}(q-1)^{2}C_{\ell}^{2}/n_{j}}\right).$
Then, we have
$\mathbb{P}\left(\mathfrak{D}^{-}(\bm{g})\right)=$ $p\left(\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})\leq\beta\right)$
$\leq$ $p\left(\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})\leq\sum\nolimits_{1\leq j<k\leq q}Mp(y_{j}=0,y_{k}=0)\right)$
$\leq$ $p\left(\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})\leq\sum\nolimits_{1\leq j<k\leq q}Mp(y_{j}=0,y_{k}=0)+R_{\mathrm{R}}^{\ell}(\bm{g})-\gamma\right)$
$\leq$ $\exp\left(\frac{-2\gamma^{2}}{\sum_{j=1}^{q}(1-\pi_{j})^{2}(q-1)^{2}C_{\ell}^{2}/n_{j}}\right),$
which concludes the proof. ∎
Then, we give the proof of Theorem 5.
Proof of Theorem 5.
To begin with, we have
$\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\sum\nolimits_{1\leq j<k\leq q}Mp(y_{j}=0,y_{k}=0)-R_{\mathrm{R}}^{\ell}(\bm{g})=\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})-\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]\geq 0.$
Besides, we have
$\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})\leq C_{\ell}(q-1)\sum\nolimits_{j=1}^{q}(1-\pi_{j}).$
Then,
$\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\sum\nolimits_{1\leq j<k\leq q}Mp(y_{j}=0,y_{k}=0)-R_{\mathrm{R}}^{\ell}(\bm{g})$
$=$ $\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})-\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]$
$=$ $\int_{\widehat{\mathcal{D}}\in\mathfrak{D}^{-}(\bm{g})}\left(\left|\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})-\beta\right|+\beta-\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})\right)p\left(\widehat{\mathcal{D}}\right)\mathop{}\!\mathrm{d}\widehat{\mathcal{D}}$
$\leq$ $\mathop{\sup}_{\widehat{\mathcal{D}}\in\mathfrak{D}^{-}(\bm{g})}\left(2\beta+2\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})\right)\int_{\widehat{\mathcal{D}}\in\mathfrak{D}^{-}(\bm{g})}p\left(\widehat{\mathcal{D}}\right)\mathop{}\!\mathrm{d}\widehat{\mathcal{D}}$
$=$ $\mathop{\sup}_{\widehat{\mathcal{D}}\in\mathfrak{D}^{-}(\bm{g})}\left(2\beta+2\hat{R}_{\mathrm{R}}^{\ell}(\bm{g})\right)\mathbb{P}\left(\mathfrak{D}^{-}(\bm{g})\right)$
$\leq$ $\left(2\beta+2C_{\ell}(q-1)\sum\nolimits_{j=1}^{q}(1-\pi_{j})\right)\exp\left(\frac{-2\gamma^{2}}{\sum_{j=1}^{q}(1-\pi_{j})^{2}(q-1)^{2}C_{\ell}^{2}/n_{j}}\right),$
which concludes the first part of the proof. Then we provide an upper bound for $\left|\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})-\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]\right|$. When an instance from $\mathcal{D}_{j}$ is replaced by another instance, the value of $\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})$ changes at most $(1-\pi_{j})(q-1)C_{\ell}/n_{j}$. Therefore, by applying the McDiarmid’s inequality, we have the following inequalities with probability at least $1-\delta/2$:
$\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})-\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]\leq$ $\sum\nolimits_{j=1}^{q}(1-\pi_{j})(q-1)C_{\ell}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n_{j}}},$
$\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\leq$ $\sum\nolimits_{j=1}^{q}(1-\pi_{j})(q-1)C_{\ell}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n_{j}}}.$
Therefore, we have the following inequalities with probability at least $1-\delta$:
$\left|\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})-\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]\right|\leq$ $\sum\nolimits_{j=1}^{q}(1-\pi_{j})(q-1)C_{\ell}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n_{j}}}.$
Finally,
$\left|\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})-\sum\nolimits_{1\leq j<k\leq q}Mp(y_{j}=0,y_{k}=0)-R_{\mathrm{R}}^{\ell}(\bm{g})\right|$
$=$ $\left|\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})-\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]+\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\sum\nolimits_{1\leq j<k\leq q}Mp(y_{j}=0,y_{k}=0)-R_{\mathrm{R}}^{\ell}(\bm{g})\right|$
$\leq$ $\left|\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})-\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]\right|+\left|\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\sum\nolimits_{1\leq j<k\leq q}Mp(y_{j}=0,y_{k}=0)-R_{\mathrm{R}}^{\ell}(\bm{g})\right|$
$\leq$ $\sum\nolimits_{j=1}^{q}(1-\pi_{j})(q-1)C_{\ell}\sqrt{\frac{\ln{\left(2/\delta\right)}}{2n_{j}}}$
$+\left(2\beta+2C_{\ell}(q-1)\sum\nolimits_{j=1}^{q}(1-\pi_{j})\right)\exp\left(\frac{-2\gamma^{2}}{\sum_{j=1}^{q}(1-\pi_{j})^{2}(q-1)^{2}C_{\ell}^{2}/n_{j}}\right),$ (27)
which concludes the proof. ∎
B.8 Proof of Theorem 6
Lemma 9.
Based on the above assumptions, for any $\delta>0$, the following inequality holds with probability at least $1-\delta$:
$\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|R_{\mathrm{R}}^{\ell}(\bm{g})+\sum\nolimits_{j<k}Mp(y_{j}=0,y_{k}=0)-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|\leq\left(2\beta+2C_{\ell}(q-1)\sum\nolimits_{j=1}^{q}(1-\pi_{j})\right)\Delta^{\prime}$
$+\sum\nolimits_{j=1}^{q}(1-\pi_{j})(q-1)C_{\ell}\sqrt{\frac{\ln{\left(1/\delta\right)}}{n_{j}}}+\sum\nolimits_{j=1}^{q}4L_{\ell}(q-1)(1-\pi_{j})\mathfrak{R}_{n_{j},p_{j}}\left(\mathcal{G}\right).$ (28)
Proof.
When an instance in $\mathcal{D}_{j}$ is replaced by another instance, the value of $\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|$ changes at most $(1-\pi_{j})(q-1)C_{\ell}/n_{j}$. Therefore, by applying the McDiarmid’s inequality, we have the following inequalities with probability at least $1-\delta$:
$\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|-\mathbb{E}\left[\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|\right]$
$\leq$ $\sum\nolimits_{j=1}^{q}(1-\pi_{j})(q-1)C_{\ell}\sqrt{\frac{\ln{\left(1/\delta\right)}}{n_{j}}}.$ (29)
Then,
$\mathbb{E}\left[\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|\right]$
$=$ $\mathbb{E}_{\widehat{\mathcal{D}}}\left[\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}_{\widehat{\mathcal{D}}^{\prime}}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|\right]$
$\leq$ $\mathbb{E}_{\widehat{\mathcal{D}},\widehat{\mathcal{D}}^{\prime}}\left[\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g};\widehat{\mathcal{D}})-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g}^{\prime};\widehat{\mathcal{D}}^{\prime})\right|\right],$ (30)
[…]
Lemma 10.
If $\ell:\mathbb{R}\times\left\{0,1\right\}\rightarrow\mathbb{R}$ is a Lipschitz continuous function with a Lipschitz constant $L_{\ell}$ and satisfies $\forall y,\ell(0,y)=0$, we have
$\mathfrak{R}^{\prime}_{n,p}\left(\ell\circ(\mathcal{G}-\mathcal{G})\right)\leq 4L_{\ell}\mathfrak{R}^{\prime}_{n,p}(\mathcal{G}),$
where $\ell\circ((\mathcal{G}-\mathcal{G}))=\{\ell\circ\left(g_{j}-g_{k}\right)|g_{j}\in\mathcal{G},g_{k}\in\mathcal{G}\}$.
Proof.
$\mathfrak{R}^{\prime}_{n,p}\left(\ell\circ(\mathcal{G}-\mathcal{G})\right)$
$=$ $2\mathfrak{R}^{\prime}_{n,p}\left(\ell\circ(\mathcal{G})\right)$
$\leq$ $4L_{\ell}\mathfrak{R}^{\prime}_{n,p}(\mathcal{G}),$
where the first inequality is by symmetrization [Mohri et al., 2012] and the second inequality is by Lemma 4. The proof is complete. ∎
Therefore, we have
$\left|\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g};\widehat{\mathcal{D}})-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g};\widehat{\mathcal{D}}^{\prime})\right|$
$=$ $\left|\left|\hat{R}_{\mathrm{R}}^{\ell}(\bm{g};\widehat{\mathcal{D}})-\beta\right|-\left|\hat{R}_{\mathrm{R}}^{\ell}(\bm{g};\widehat{\mathcal{D}}^{\prime})-\beta\right|\right|$
$\leq$ $\left|\hat{R}_{\mathrm{R}}^{\ell}(\bm{g};\widehat{\mathcal{D}})-\hat{R}_{\mathrm{R}}^{\ell}(\bm{g};\widehat{\mathcal{D}}^{\prime})\right|$
$=$ $\left|\sum\nolimits_{1\leq j<k\leq q}\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\left(\ell\left(g_{j}(\bm{x}_{i}^{j})-g_{k}(\bm{x}_{i}^{j}),0\right)-\ell\left(g_{j}(\bm{x}_{i}^{j^{\prime}})-g_{k}(\bm{x}_{i}^{j^{\prime}}),0\right)\right)\right$.
$\left.+\frac{1-\pi_{k}}{n_{k}}\sum\nolimits_{i=1}^{n_{k}}\left(\ell\left(g_{j}(\bm{x}_{i}^{k})-g_{k}(\bm{x}_{i}^{k}),1\right)-\ell\left(g_{j}(\bm{x}_{i}^{k^{\prime}})-g_{k}(\bm{x}_{i}^{k^{\prime}}),1\right)\right)\right|$
$\leq$ $\sum\nolimits_{1\leq j<k\leq q}\left|\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\left(\ell\left(g_{j}(\bm{x}_{i}^{j})-g_{k}(\bm{x}_{i}^{j}),0\right)-\ell\left(g_{j}(\bm{x}_{i}^{j^{\prime}})-g_{k}(\bm{x}_{i}^{j^{\prime}}),0\right)\right)\right|$
$+\sum\nolimits_{1\leq j<k\leq q}\left|\frac{1-\pi_{k}}{n_{k}}\sum\nolimits_{i=1}^{n_{k}}\left(\ell\left(g_{j}(\bm{x}_{i}^{k})-g_{k}(\bm{x}_{i}^{k}),1\right)-\ell\left(g_{j}(\bm{x}_{i}^{k^{\prime}})-g_{k}(\bm{x}_{i}^{k^{\prime}}),1\right)\right)\right|$
$=$ $\sum\nolimits_{1\leq j<k\leq q}\left|\frac{1-\pi_{j}}{n_{j}}\sum\nolimits_{i=1}^{n_{j}}\left(\bar{\ell}\left(g_{j}(\bm{x}_{i}^{j})-g_{k}(\bm{x}_{i}^{j}),0\right)-\bar{\ell}\left(g_{j}(\bm{x}_{i}^{j^{\prime}})-g_{k}(\bm{x}_{i}^{j^{\prime}}),0\right)\right)\right|$
$+\sum\nolimits_{1\leq j<k\leq q}\left|\frac{1-\pi_{k}}{n_{k}}\sum\nolimits_{i=1}^{n_{k}}\left(\bar{\ell}\left(g_{j}(\bm{x}_{i}^{k})-g_{k}(\bm{x}_{i}^{k}),1\right)-\bar{\ell}\left(g_{j}(\bm{x}_{i}^{k^{\prime}})-g_{k}(\bm{x}_{i}^{k^{\prime}}),1\right)\right)\right|,$ (31)
where the inequalities are due to the triangle inequality. Then, by combining Inequalities 30 and 31, it is a routine work [Mohri et al., 2012] to show that
$\mathbb{E}\left[\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|\right]$
$\leq$ $\sum\nolimits_{j=1}^{q}(q-1)(1-\pi_{j})\mathfrak{R}^{\prime}_{n_{j},p_{j}}\left(\bar{\ell}\circ(\mathcal{G}-\mathcal{G})\right)$
$\leq$ $\sum\nolimits_{j=1}^{q}4L_{\ell}(q-1)(1-\pi_{j})\mathfrak{R}^{\prime}_{n_{j},p_{j}}\left(\mathcal{G}\right)$
$=$ $\sum\nolimits_{j=1}^{q}4L_{\ell}(q-1)(1-\pi_{j})\mathfrak{R}_{n_{j},p_{j}}\left(\mathcal{G}\right),$ (32)
where the second inequality is by Lemma 10 and the last equality is by Lemma 3. By combining Inequalities B.8 and 32, we have the following inequalities with probability at least $1-\delta$:
$\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|$
$\leq$ $\sum\nolimits_{j=1}^{q}(1-\pi_{j})(q-1)C_{\ell}\sqrt{\frac{\ln{\left(1/\delta\right)}}{n_{j}}}+\sum\nolimits_{j=1}^{q}4L_{\ell}(q-1)(1-\pi_{j})\mathfrak{R}_{n_{j},p_{j}}\left(\mathcal{G}\right).$ (33)
Finally, we have the following inequality with probability at least $1-\delta$:
$\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|R_{\mathrm{R}}^{\ell}(\bm{g})+\sum\nolimits_{j<k}Mp(y_{j}=0,y_{k}=0)-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|$
$=$ $\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|R_{\mathrm{R}}^{\ell}(\bm{g})+\sum\nolimits_{j<k}Mp(y_{j}=0,y_{k}=0)-\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]+\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|$
$\leq$ $\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|R_{\mathrm{R}}^{\ell}(\bm{g})+\sum\nolimits_{j<k}Mp(y_{j}=0,y_{k}=0)-\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]\right|+\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|\mathbb{E}\left[\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right]-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|$
$\leq$ $\left(2\beta+2C_{\ell}(q-1)\sum\nolimits_{j=1}^{q}(1-\pi_{j})\right)\Delta^{\prime}+\sum\nolimits_{j=1}^{q}(1-\pi_{j})(q-1)C_{\ell}\sqrt{\frac{\ln{\left(1/\delta\right)}}{n_{j}}}$
$+\sum\nolimits_{j=1}^{q}4L_{\ell}(q-1)(1-\pi_{j})\mathfrak{R}_{n_{j},p_{j}}\left(\mathcal{G}\right),$
where the last inequality is by Inequalities B.8 and 15. The proof is complete. ∎
Then, we provide the proof of Theorem 6.
Proof of Theorem 6.
$R_{\mathrm{R}}^{\ell}(\tilde{\bm{g}}_{\mathrm{R}})-R_{\mathrm{R}}^{\ell}(\bm{g}^{*}_{\mathrm{R}})$
$=$ $R_{\mathrm{R}}^{\ell}(\tilde{\bm{g}}_{\mathrm{R}})+\sum\nolimits_{j<k}Mp(y_{j}=0,y_{k}=0)-\tilde{R}_{\mathrm{R}}^{\ell}(\tilde{\bm{g}}_{\mathrm{R}})+\tilde{R}_{\mathrm{R}}^{\ell}(\tilde{\bm{g}}_{\mathrm{R}})-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g}^{*}_{\mathrm{R}})+\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g}^{*}_{\mathrm{R}})$
$-\sum\nolimits_{j<k}Mp(y_{j}=0,y_{k}=0)-R_{\mathrm{R}}^{\ell}(\bm{g}^{*}_{\mathrm{R}})$
$\leq$ $R_{\mathrm{R}}^{\ell}(\tilde{\bm{g}}_{\mathrm{R}})+\sum\nolimits_{j<k}Mp(y_{j}=0,y_{k}=0)-\tilde{R}_{\mathrm{R}}^{\ell}(\tilde{\bm{g}}_{\mathrm{R}})+\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g}^{*}_{\mathrm{R}})-\sum\nolimits_{j<k}Mp(y_{j}=0,y_{k}=0)-R_{\mathrm{R}}^{\ell}(\bm{g}^{*}_{\mathrm{R}})$
$\leq$ $2\mathop{\sup}_{g_{1},g_{2},\ldots,g_{q}\in\mathcal{G}}\left|R_{\mathrm{R}}^{\ell}(\bm{g})+\sum\nolimits_{j<k}Mp(y_{j}=0,y_{k}=0)-\tilde{R}_{\mathrm{R}}^{\ell}(\bm{g})\right|.$
Then, based on Lemma 9, the proof is complete. ∎
B.9 Proof of Corollary 2
Lemma 11 (Theorem 10 in Gao and Zhou [2013]).
If $\ell$ is a differentiable and non-increasing function such that $\forall y,\ell^{\prime}(0,y)<0$ and $\ell(z,y)+\ell(-z,y)=M$, then Eq. (4) is consistent w.r.t. the ranking loss.
Then we provide the proof of Corollary 2.
Proof of Corollary 2.
Since the proposed risk in Eq. (1) is equivalent to the risk in Eq. (4), it is sufficient to prove that for any sequence $\left\{\bm{g}_{t}\right\}$ that if $R^{\ell}_{\mathrm{R}}(\bm{g}_{t})\rightarrow R^{\ell*}_{\mathrm{R}}$, then $R^{\mathrm{0-1}}_{\mathrm{R}}(\bm{f}_{t})\rightarrow R^{*}_{\mathrm{R}}$. ∎
Appendix C Details of Experiments
C.1 Synthetic Dataset Generation
For synthetic datasets, we generated partial labels in two ways. In Case-a, we flipped negative labels to positive labels independently with probability $\tau$. In Case-b, for each class we deterministically enforced a flip rate of $\rho$ by uniformly sampling a $\rho$ fraction of negative instances in that class and flipping them to positive. We set $\tau{=}\rho{=}0.9$ in this paper.
C.2 Baseline
We evaluate against five classical baselines commonly used in PML/CML learning. (A) BCE: uses the given partial label as the cross-entropy target. (B) CCMN [Xie and Huang, 2023]: treats PML as multi-label classification with class-conditional noise, relying on a noise transition matrix. (C) GDF [Gao et al., 2023]: proposes an unbiased risk estimator for multi-labeled single complementary label learning. (D) CTL [Gao et al., 2025]: introduces a risk-consistent approach by rewriting the loss function. (E) MLCL [Gao et al., 2024]: estimates an initial transition matrix via binary decompositions, then refines it with label correlations.
C.3 Implementation Details
For real-world datasets, we used an MLP encoder for all baselines, trained for 200 epochs with a learning rate of 5e-3, weight decay of 1e-4, and the SGD optimizer with cosine decay. For synthetic image datasets, we adopted a ResNet-50 backbone pretrained on ImageNet [Deng et al., 2009], trained for 30 epochs with a learning rate of 1e-4 using the Adam optimizer. We used the binary cross-entropy loss. For fair comparisons, we used the same setup across all baselines. We assumed that the class priors were accessible to the learning algorithm.